The equilibrium constant for the following reaction:

H2(g) + Br2(g) ↔ 2HBr (g)

is 1.20 x 103 at a certain temperature. Find the equilibrium pressure of HBr if 9.20 atm of HBr is introduced into a sealed container at this temperature.
Answer #1

2 HBr ig, F H₂ 1gg + Br2 lg.) - 9.2 atmo os (9.2 - ~) (072) 10/22 (At t=0) (At t=t) Kp = [PH2 ] [ Pers] [Per]? («/2) (0/2) (9Hence is Equilibrium pressure of HBr Puer = (4.2-4) THBg = 9,2 - 0.5312 PHBr = 8.6688 atm